Using yellow, cyan, and magenta as primary colors, express green, red, and blue in the vector form.
Answer #1

Since, if Red, Blue and Green are primary colours, the relation is as follows

R+G=Y...........(i)

R+B=M..........(ii)

G+B=C..........(iii)

Y=YELLOW

M=MAGENTA

C=CYAN

If we subtract equation (ii) from (i) we will get

G-B=Y-M.......(iv)

Now adding (iii) and (iv), we get

G+B+G-B=C+Y-M

2G=C+Y-M

Hence

On substituting the value of G in (i), we get

R=Y-G

On substituting the value R in (ii) we get

B=M-R